OpenStudy (anonymous):

Determine the slope of the line graphed by 5y + 6x - 2 = 0. Type a numerical answer in the space provided. If necessary, use the / key to represent a fraction bar and leave your answer in terms of an improper fraction.

jimthompson5910 (jim_thompson5910):

what do you get when you solve for y?

OpenStudy (anonymous):

You just have to solve for y and see the value of the constant that is multiplying x.

OpenStudy (anonymous):

huh ?

jimthompson5910 (jim_thompson5910):

let's completely isolate y 5y + 6x - 2 = 0 5y + 6x = 2 5y = 2 - 6x 5y = -6x + 2 y = (-6x + 2)/5 y = (-6x)/5 + 2/5 y = (-6/5)x + 2/5 the equation is now in y = mx+b form where m = -6/5 and b = 2/5